A ball is thrown vertically upward with a speed of 15 m/s. Determine the time of flight when it

returns to its original position.


1
Expert's answer
2021-10-01T12:18:43-0400

Explanations & Calculations


  • Since it returns to the original position, the displacement is zero. Therefore, using the equation"\small S=\small ut+\frac{1}{2}at^2" for the ball's motion airborne time can be found.

"\qquad\qquad
\begin{aligned}
\small 0&=\small 15\times t+\frac{1}{2}\times(-g)\times t^2 \\
&=\small 15t-4.9t^2 \\
\small 0&=\small t\Big(15-4.9t\Big)\\
\\
\small t &=\begin{cases}
\small 0:extraneouse\\
 \frac{15}{4.9}=\small 3.1: accepted
\end{cases}
\end{aligned}"

  • Therefore, the time of flight is "\small \bold{3.1\,s}"